OpenStudy (anonymous):

Which ordered pair is a solution of the equation 5x + y = –23?

OpenStudy (anonymous):

Plug in the numbers and see if they satisfy the equation. 5(-3) + (-4) = -23 -15 - 4 = -23 -19 = -23 Nope! That one doesn't do it. Now you try the other choices.

OpenStudy (anonymous):

Which ordered pair is a solution of the equation 5x + y = –23? Answers (–3, –4) (–5, –3) (–4, –3) (–3, –5)

OpenStudy (anonymous):

based on the answer i just posted, which ordered pair makes sense to you?

OpenStudy (anonymous):

I'm confused..

OpenStudy (anonymous):

your equation 5(x)+y=-23 plug in the numbers for x and y and you have your ordered pairs.

OpenStudy (anonymous):

(-4,-3) is the answer then

OpenStudy (anonymous):

there you go!

OpenStudy (anonymous):

Thanks..